A balloon ascends vertically with a constant speed for 5 sec. When a pebble falls from it, reaches the ground in 5 sec. The speed of the balloon is?

Grade:12th pass

2 Answers

Vivek Kumar Singh
6 years ago
balloon ascends vertically with a constant speed for 5 sec. then pebble Falls from it reaches the ground in 5 sec.TIME OF ASCEND=TIME OF DESCEND= u/g= 5sec.u/10=5u=50 m/sec.THE SPEED OF BALLOON IS= 50 m/sec.
AMARPREET KAUR LOTTE
4 years ago
let velocity of balloon be v so for 5seconds its distance travelled is 5v.(distance =v*t)
Now when pebble will fall from balloon it will acquire velocity of balloon in opposite direction -v and will cover same distance 5v for 5 seconds 
by applying S=UT + ½ GT^2  putting the values 
U=-V ,T=5SEC ,G=10, S=5V
5v= -5v + ½ *10*5 ^2 
5v=-5v+125
10v=125
v=12.5 
answer